OpenStudy (anonymous):

solve the linear equation 15-y=7

OpenStudy (gudden):

15 - 7 = y it means y = 8

OpenStudy (secret66):

Explain it ^_^

HanAkoSolo (jamierox4ev3r):

@MitziYata welcome to OpenStudy! Ask as many questions as you want, and feel free to talk to me if you need anything! Do you understand the problem? If not, I'll explain it to you @gudden , careful here. At OpenStudy, we don't want to provide answers without helping the user understand them! please keep this in mind for future questions. Thank you

OpenStudy (gudden):

haha @secret66 .... 15 - y = 7 now bring the y to the LHS nad 7 to the RHS and u would get 15 - 7 = y now i hope... u atleast know... how to subtract...:D

OpenStudy (gudden):

@Jamierox4ev3r .. alright.. next time i'll be careful about it..

HanAkoSolo (jamierox4ev3r):

hey @gudden nice, you redeemed yourself! Now I know it may seem tedious, but its important to show this process so that the user understands! Plus, its in the rules >_> To break it down visibly: 15-y=7 -15 -15 ------- y=7-15

OpenStudy (secret66):

xD I just told you to do that because maybe the user don't know how you solved it.. And ask the same question again and again..So its better to explain it than to give direct answers

HanAkoSolo (jamierox4ev3r):

ditto that ^

OpenStudy (secret66):

LOL Jamie

HanAkoSolo (jamierox4ev3r):

well that's practically what i wanted to say xDD so congrats ^_^

HanAkoSolo (jamierox4ev3r):

Now we're all buddies…but i wonder if the original asker of the question understand the answer…. LOL

OpenStudy (gudden):

hahah...!... okay guys... @Jamierox4ev3r u forgot to put the minus sign before y... in the last line.... and sure... i would explain the same way next time...:D

OpenStudy (gudden):

as u have explained me :D

HanAkoSolo (jamierox4ev3r):

oops xDD but yeah, I'm glad we all came to a mutual agreement.

OpenStudy (gudden):

ofcourse..! am glad too...:D

OpenStudy (secret66):

*bows* Imfao

OpenStudy (anonymous):

Thank you for the answer! I understand it well now, my teacher gave me linear equations as homework and she didn't even teach us how to do it at all :/

OpenStudy (gudden):

oh... sweety... u could have asked me... if u didn't understand it then....!! and am even sorry for not explaining u that easy.. cuz i belong to India.. and here... the linear is taught to us along with the 2 variables...so.. sorry sweety..!.. cuz i explained u the way were taught... anyways im gald that u understood it welll

OpenStudy (secret66):

XD She lives in India so ppl living in India give direct answers..Not explanation hehe! JK

OpenStudy (anonymous):

Haha don't worry it's alright, I understood it fine :) thank you again!

OpenStudy (gudden):

hahah,...!. yeah..! we don't believe in wasting our tym in explanations...:p ( except in exams..:p... cuz then every sentence we write has its own value.....)..:D